OpenStudy (anonymous):

A baseball player swings and hits a pop fly straight up in air to the catcher. the height of the baseball in meters t seconds after it is hit is given by the quadratic function h(t)=-4.9t^(2)+34.3t+1. How long does it take for the baseball to reach its maximum height? What is the maximum height obtained by the baseball?

OpenStudy (anonymous):

t=-b/2a is the time to reach it's maximum height. So t=-34.3/(2)(-4.9) Which is 3.5 seconds. Plug in 3.5 seconds into your equation to get the height. So you have\[h(t)=-4.9(3.5)^{2}+34.3(3.5)+1\] Which is about 61.03 meters